# MCAQ-YOLO: Morphological Complexity-Aware Quantization for Efficient Object Detection with Curriculum Learning

### v1

@@ IV-E Training Objective
Detection loss : standard YOLO loss with classification, localization, and objectness terms.
@@
Knowledge distillation loss : aligns the quantized model with an FP32 teacher using both logit and feature-level matching.
Regularization : penalty on mapping network weights.
@@ V-A1 Datasets
Construction Safety Equipment (CSE) (primary): 12,000 images (10,000 train, 1,000 validation, 1,000 test) with six classes (Person, Helmet, Vest, Gloves, Boots, Goggles) and 58,342 bounding boxes. Images are collected from multiple publicly available PPE datasets and additional in-house images. The dataset exhibits substantial variation in background clutter, viewpoint, and object scale.
Standard benchmarks: To assess generalization, we evaluate on a 10,000-image subset of COCO 2017, Pascal VOC 2012, and Cityscapes. For COCO, we focus on 20 representative categories to maintain manageable evaluation time under mixed-precision configurations.
V-A2 Implementation Details
All models are trained in PyTorch 2.1.0 with CUDA 12.0. Training uses an AdamW optimizer (weight decay 0.05, , ), cosine learning rate schedule with warmup, batch size 64 (16 per GPU with gradient accumulation), and a total of 300 epochs (warmup, standard training, and fine-tuning phases). Unless stated otherwise, we report results on YOLOv8 with 4-bit average precision under TensorRT 8.6 deployment.
V-A3 Evaluation Metrics
We report mAP@0.5, mAP@[0.5:0.95], AP for small/medium/large objects, AR@100, compressed model size, FPS, and energy per image. Each configuration is run with 10 random seeds, and 95% confidence intervals are estimated via bootstrap resampling.
@@ V-B Main Results on CSE Dataset
Method Bits mAP (%) Efficiency W A @.5 @[.5:.95] Small Med Large AR@100 Size (MB) FPS Energy (J) YOLOv8-FP32 32 32 89.30.3 68.10.4 42.30.5 71.20.4 84.60.3 72.40.3 108.3 92 0.45 YOLOv8-FP16 16 16 89.10.3 67.90.4 42.00.5 71.00.4 84.30.3 72.10.3 54.2 118 0.32 Uniform-8bit 8 8 88.10.3 66.90.3 40.80.4 70.10.4 83.20.3 71.20.3 27.1 134 0.21 Uniform-4bit 4 4 82.10.4 58.30.5 31.20.6 62.40.5 76.80.4 63.50.4 13.5 156 0.14 Uniform-3bit 3 3 74.30.6 48.20.7 21.30.8 51.30.6 68.40.5 53.20.5 10.1 168 0.11 LSQ [10] 4 4 83.20.3 59.80.4 32.60.5 63.70.4 77.90.3 64.80.3 13.5 156 0.14 PACT [11] 4 4 82.80.4 59.10.5 32.10.5 63.20.5 77.40.4 64.20.4 13.5 156 0.14 DSQ [12] 4 4 83.00.4 59.50.4 32.40.5 63.50.4 77.70.4 64.50.4 13.5 155 0.14 HAQ [13] mix mix 84.10.3 61.20.4 34.30.5 65.10.4 78.90.3 66.20.3 15.2 148 0.17 HAWQ-V3 [4] mix mix 84.70.3 62.10.4 35.10.5 65.90.4 79.60.3 67.00.3 15.8 145 0.18 Reducing Osc. [5] 4 4 83.90.3 60.50.4 33.20.4 64.30.4 78.50.3 65.40.3 13.5 154 0.14 Info. Entropy [6] mix mix 84.30.4 61.60.4 34.70.5 65.50.4 79.20.4 66.50.4 15.5 146 0.17 MCAQ (Ours) 4.2 4.2 85.60.4 63.10.4 36.30.5 67.00.4 80.70.4 68.20.4 14.2 151∗ 0.14 MCAQ-Fast 4.0 4.0 84.80.4 62.30.4 35.40.5 66.20.4 79.90.4 67.40.4 13.5 158∗ 0.14
On the CSE dataset, MCAQ-YOLO with an average of 4.2 bits achieves 85.6% mAP@0.5, outperforming uniform 4-bit quantization by 3.5 percentage points and the strongest mixed-precision baseline (HAWQ-V3) by 0.9 percentage points, while maintaining a 7.63 compression ratio (108.3 MB → 14.2 MB) and comparable runtime. Notably, MCAQ achieves these gains with lower average bit-width (4.2 bits) compared to HAWQ-V3’s mixed-precision scheme. The improvements are particularly pronounced for small objects (36.3% vs HAWQ-V3’s 35.1%), which tend to be morphologically complex.
@@ V-C Cross-Dataset Generalization
Table III reports results on COCO 2017 and Pascal VOC 2012. The performance gains compared to uniform 4-bit quantization vary across datasets: CSE (+3.5%), COCO (+2.9%), and VOC (+2.3%). When compared to the strongest baseline (HAWQ-V3), MCAQ still shows consistent improvements: CSE (+0.9%), COCO (+1.3%), and VOC (+0.9%). This variation correlates with dataset characteristics—CSE contains more morphologically diverse objects (construction workers with varying poses and equipment), while COCO and VOC have more uniform object distributions. The consistent positive gains across all datasets suggest that morphological complexity provides a useful but dataset-dependent signal for spatial quantization.
Method COCO 2017 Pascal VOC 2012 mAP@.5 mAP@[.5:.95] APS FPS mAP@.5 mAP@[.5:.95] APS FPS YOLOv8-FP32 51.2 32.4 15.3 95 82.3 52.1 38.2 98 Uniform-4bit 45.2 28.1 11.2 162 76.8 48.3 33.4 168 HAWQ-V3 46.8 29.5 12.4 152 78.2 49.8 34.7 156 Info. Entropy 47.1 29.7 12.6 150 78.5 49.9 35.0 154 MCAQ 48.1 30.3 13.2 158 79.1 50.4 35.8 163
@@ V-D Morphological Complexity and Quantization Sensitivity
Table IV illustrates average morphological complexity and mAP degradation under 3-bit quantization for representative classes. The Spearman correlation between complexity and mAP drop is 0.89 (p 0.001), supporting the core hypothesis.
Class Bits mAP Drop Person 1.68.06 0.82.05 0.71.04 0.38.03 6.2.4 0.72.03 5.8.2 17.2% Helmet 1.25.04 0.33.03 0.30.02 0.10.01 1.8.2 0.25.02 4.1.1 5.3% Background 1.15.03 0.28.02 0.25.02 0.08.01 1.2.1 0.21.01 3.8.1 2.1% Spearman correlation: vs. mAP drop
V-E Runtime Analysis and Ablation Studies
A detailed runtime breakdown is reported in Table V. With all optimizations enabled, morphological analysis and bit allocation together add approximately 1.8 ms per image (1.4 ms for morphological analysis + 0.4 ms for bit allocation and quantization), demonstrating that the method introduces minimal runtime overhead while achieving significant accuracy gains.
Component CPU (ms) GPU (ms) Optimized (ms) Peak Mem (MB) Avg Mem (MB) YOLOv8 Backbone 15.2 4.8 4.8 45.2 32.1 Morphological Analysis 12.8 3.5 1.4 31.7 22.0 Bit Allocation + Quant. 3.6 1.0 0.4 3.3 2.3 Total Overhead 16.4 4.5 1.8 35.0 24.3 Total with YOLO 31.6 9.3 6.6 76.9 54.1
Ablation studies (Table VI) show that fractal dimension and feature interaction terms contribute the most to performance, whereas curriculum learning provides clear but moderate gains in both convergence speed and final accuracy.
Configuration mAP@0.5 FPS Full MCAQ 85.2 – 151 Component Ablation w/o Fractal dimension 84.4 0.8 155 w/o Texture entropy 84.7 0.5 154 w/o Gradient variance 84.9 0.3 153 w/o Edge density 84.8 0.4 154 w/o Contour complexity 84.6 0.6 155 w/o Interaction terms 84.0 1.2 152 w/o Curriculum 84.3 0.9 151 Tile Size Variations Tile size = 16 85.4 0.2 143 Tile size = 32 (default) 85.2 – 151 Tile size = 64 84.6 0.6 158 Complexity Threshold = 0.4 84.5 0.7 150 = 0.6 (default) 85.2 – 151 = 0.8 84.7 0.5 152 Bit Range 3-5 bits 84.1 1.1 156 3-6 bits (default) 85.2 – 151 4-6 bits 85.0 0.2 148 Linear mapping only 83.8 1.4 153 Layer-wise only (no spatial) 83.7 1.5 155
V-F Convergence Behavior
@@
Figure 1 illustrates training curves with and without curriculum learning, as well as a standard QAT baseline. The curriculum-based strategy reaches 95% of its final performance (81.9% mAP) at 20k iterations, while the non-curriculum variant requires 50k iterations to reach a comparable level (81.1% mAP), demonstrating 2.5× faster convergence. The curriculum approach also achieves 3.2 percentage points higher final accuracy (85.6% vs 82.4%), showing improvements in both training efficiency and model performance.
V-G Qualitative Analysis
Figure 2 shows qualitative examples in which MCAQ-YOLO assigns higher bit-widths to complex regions (e.g., persons and articulated objects) and lower bit-widths to relatively uniform backgrounds, while maintaining accurate detections.
@@ VI-A Key Findings and Implications
Curriculum learning proves to be an effective mechanism for stabilizing quantization-aware training. By gradually increasing the complexity of training samples and adjusting the temperature of bit allocation, we obtain faster convergence, lower gradient variance, and slightly improved final accuracy.

### v2

@@ IV-E Training Objective
Detection loss : Standard YOLO loss comprising classification, localization (CIoU), and objectness terms.
@@
Knowledge distillation loss : Aligns the quantized model with an FP32 teacher using both logit-level and feature-level matching.
Regularization : penalty on mapping network weights with coefficient .
The time-varying coefficient is annealed from 0.01 to 0.1 during training to gradually enforce the bit budget constraint.
@@ V-A1 Datasets
Construction Safety Equipment (CSE) (primary benchmark): 12,000 images (10,000 train, 1,000 validation, 1,000 test) with six classes (Person, Helmet, Vest, Gloves, Boots, Goggles) and 58,207 bounding boxes after filtering invalid annotations. Sources, cleaning protocol, and split methodology are detailed in Appendix D; we ensure no near-duplicate images across splits via perceptual hashing. The dataset exhibits substantial variation in background clutter, viewpoint, and object scale, making it suitable for evaluating morphology-driven quantization.
Standard benchmarks: To assess generalization, we evaluate on COCO 2017 val (5,000 images, 80 classes) and Pascal VOC 2012 (5,717 images, 20 classes).
V-A2 Implementation and Evaluation
Standard YOLOv8 [30] training and TensorRT 8.6 deployment protocols are followed; full hyperparameters, energy measurement settings, and statistical analysis procedures are provided in Appendices F–C. Baseline fairness: All methods are deployed under the same TensorRT 8.6 pipeline with identical input preprocessing and batch size; baseline implementations (LSQ, PACT, HAWQ-V3, etc.) use their official codebases without modification, while only MCAQ employs the tile-wise quantization kernel. The reported metrics include mAP@0.5, mAP@[0.5:0.95], AP for small/medium/large objects, AR@100, compressed model size, FPS, and energy per image. Each configuration is run with 10 random seeds, and 95% confidence intervals are estimated via bootstrap resampling (10,000 samples). Statistical significance is assessed using paired -tests with Benjamini–Hochberg correction for multiple comparisons.
@@ V-B Main Results on CSE Dataset
| Method | Bits | mAP (%) | Efficiency | ||||||||
| W | A | @.5 | @[.5:.95] | Small | Med | Large | AR@100 | Size | FPS | Energy | |
| YOLOv8-FP32 | 32 | 32 | 89.30.3 | 68.10.4 | 42.30.5 | 71.20.4 | 84.60.3 | 72.40.3 | 108.3 | 92 | 0.45 |
| YOLOv8-FP16 | 16 | 16 | 89.10.3 | 67.90.4 | 42.00.5 | 71.00.4 | 84.30.3 | 72.10.3 | 54.2 | 118 | 0.32 |
| Uniform-8bit | 8 | 8 | 88.10.3 | 66.90.3 | 40.80.4 | 70.10.4 | 83.20.3 | 71.20.3 | 27.1 | 134 | 0.21 |
| Uniform-4bit | 4 | 4 | 82.10.4 | 58.30.5 | 31.20.6 | 62.40.5 | 76.80.4 | 63.50.4 | 13.5 | 156 | 0.14 |
| Uniform-3bit | 3 | 3 | 74.30.6 | 48.20.7 | 21.30.8 | 51.30.6 | 68.40.5 | 53.20.5 | 10.1 | 168 | 0.11 |
| LSQ [11] | 4 | 4 | 83.20.3 | 59.80.4 | 32.60.5 | 63.70.4 | 77.90.3 | 64.80.3 | 13.5 | 156 | 0.14 |
| PACT [12] | 4 | 4 | 82.80.4 | 59.10.5 | 32.10.5 | 63.20.5 | 77.40.4 | 64.20.4 | 13.5 | 156 | 0.14 |
| DSQ [13] | 4 | 4 | 83.00.4 | 59.50.4 | 32.40.5 | 63.50.4 | 77.70.4 | 64.50.4 | 13.5 | 155 | 0.14 |
| GPTQ [7] | 4 | 4 | 83.50.3 | 60.10.4 | 33.00.5 | 64.00.4 | 78.20.3 | 65.10.3 | 13.5 | 156 | 0.14 |
| HAQ [14] | mix | mix | 84.10.3 | 61.20.4 | 34.30.5 | 65.10.4 | 78.90.3 | 66.20.3 | 15.2 | 148 | 0.17 |
| HAWQ-V3 [4] | mix | mix | 84.70.3 | 62.10.4 | 35.10.5 | 65.90.4 | 79.60.3 | 67.00.3 | 15.8 | 145 | 0.18 |
| Reducing Osc. [5] | 4 | 4 | 83.90.3 | 60.50.4 | 33.20.4 | 64.30.4 | 78.50.3 | 65.40.3 | 13.5 | 154 | 0.14 |
| Info. Entropy [6] | mix | mix | 84.30.4 | 61.60.4 | 34.70.5 | 65.50.4 | 79.20.4 | 66.50.4 | 15.5 | 146 | 0.17 |
| MCAQ (Ours) | 4 | 4.2† | 85.60.4 | 63.10.4 | 36.30.5 | 67.00.4 | 80.70.4 | 68.20.4 | 14.2 | 151∗ | 0.14 |
| MCAQ-Fast | 4 | 4.0† | 84.80.4 | 62.30.4 | 35.40.5 | 66.20.4 | 79.90.4 | 67.40.4 | 13.5 | 158∗ | 0.14 |
†Activation bits: spatial average across tiles (range 2–8 bits per tile). Weights use uniform 4-bit quantization. Size in MB, Energy in J.
On the CSE dataset, MCAQ-YOLO with uniform 4-bit weights and spatially adaptive activations (average 4.2 bits, range 2–8 per tile) achieves 85.6% mAP@0.5, outperforming uniform 4-bit quantization by 3.5 percentage points (paired -test, , ) and the strongest mixed-precision baseline (HAWQ-V3) by 0.9 percentage points, while maintaining a compression ratio (108.3 MB 14.2 MB). The improvement is particularly pronounced for small objects (+5.1% APS vs. uniform 4-bit), where precise boundary localization is critical. For latency-critical deployments, MCAQ-Fast (uniform 4-bit activations, 158 FPS) exceeds uniform 4-bit speed (156 FPS) while retaining a 2.7 percentage point mAP advantage through optimized tile boundaries.
@@ V-C Cross-Dataset Generalization
Table III reports results on COCO 2017 val and Pascal VOC 2012 to evaluate generalization beyond the primary benchmark.
| Method | COCO 2017 val | Pascal VOC 2012 | ||||||
| @.5 | @[.5:.95] | APS | FPS | @.5 | @[.5:.95] | APS | FPS | |
| YOLOv8-FP32 | 51.2 | 32.4 | 15.3 | 95 | 82.3 | 52.1 | 38.2 | 98 |
| Uniform-4bit | 45.2 | 28.1 | 11.2 | 162 | 76.8 | 48.3 | 33.4 | 168 |
| HAWQ-V3 | 46.8 | 29.5 | 12.4 | 152 | 78.2 | 49.8 | 34.7 | 156 |
| Info. Entropy | 47.1 | 29.7 | 12.6 | 150 | 78.5 | 49.9 | 35.0 | 154 |
| MCAQ | 48.1 | 30.3 | 13.2 | 158 | 79.1 | 50.4 | 35.8 | 163 |
The performance gains compared to uniform 4-bit quantization vary across datasets: CSE (+3.5%), COCO (+2.9%), and VOC (+2.3%). This variation correlates with dataset characteristics: CSE exhibits the highest morphological variability (high intra-class variation for Person), while VOC contains more uniform object appearances. Notably, the gains on standard benchmarks are modest compared to CSE—this reflects a key insight: spatial mixed-precision is most beneficial when datasets exhibit high within-image complexity variation, which is common in real-world deployment scenarios (e.g., industrial inspection, surveillance) but less pronounced in curated academic benchmarks. The primary contribution here is not marginal accuracy gains on COCO/VOC, but rather demonstrating that tile-wise spatial quantization is practically deployable in real-time detectors with minimal overhead.
@@ V-D Morphological Complexity and Quantization Sensitivity
Table IV provides a class-level analysis connecting morphological complexity to quantization vulnerability.
| Class | Bits | mAP | ||||||
| Person | 1.68 | 0.82 | 0.71 | 0.38 | 6.2 | 0.72 | 5.8 | 17.2% |
| Helmet | 1.25 | 0.33 | 0.30 | 0.10 | 1.8 | 0.25 | 4.1 | 5.3% |
| Backgr. | 1.15 | 0.28 | 0.25 | 0.08 | 1.2 | 0.21 | 3.8 | 2.1% |
Tile-level Spearman correlation (=5,000): vs. mAP drop, =0.73, <0.001. All values shown as mean (std. dev. 0.01–0.06 omitted for space).
Morphologically complex instances (e.g., Person with articulated poses and diverse textures) exhibit consistently larger mAP degradation when bit-width is reduced, while rigid and visually simple categories (e.g., Helmet and background tiles) demonstrate greater robustness under low precision. The Spearman correlation between the unified complexity score and mAP drop, computed at the tile level ( tiles), is 0.73 (), substantially higher than the correlation with activation entropy alone (). These results validate the hypothesis that morphological complexity serves as a strong predictor of quantization sensitivity.
V-E Ablation Studies
Table V presents ablation results quantifying the contribution of each component.
| Configuration | mAP@0.5 | |
| Full MCAQ-YOLO | 85.6 | – |
| w/o Curriculum learning | 83.1‡ | |
| w/o Smoothness regularization | 84.2‡ | |
| Linear mapping (no MLP) | 83.8‡ | |
| w/o Fractal + Contour | 84.9† | |
| w/o Texture + Edge | 84.1‡ | |
| Single metric only (Entropy) | 83.4‡ |
Curriculum learning contributes the largest individual gain (+2.5 pp), demonstrating that progressive exposure to harder samples stabilizes optimization under mixed precision. MLP-based mapping (+1.8 pp) outperforms linear mapping by capturing nonlinear relationships between complexity and optimal bit-width. Texture/edge metrics (+1.5 pp) prove more critical than geometric features (+0.7 pp) for the CSE dataset, likely because texture patterns dominate safety equipment appearance. Using only activation entropy yields 83.4% mAP, confirming that the proposed multi-metric approach provides meaningful improvement over entropy-only baselines. Extended ablation results are provided in Appendix B.
V-F Runtime Analysis
Deployment scenarios and overhead. MCAQ-YOLO supports two deployment modes:
(1) Static deployment (reported in Table II): For fixed test sets or known deployment environments, morphological analysis (1.8 ms) is performed once per image during calibration to precompute bit-maps, which are stored (0.5 KB per image for grid). At inference, only bit-map lookup and tile-wise quantization are required, adding 0.3 ms (156151 FPS). This mode is suitable for surveillance with fixed camera views, quality inspection with known product types, or any scenario where representative images can be pre-analyzed.
(2) Online deployment: For novel images, the full morphological analysis runs per frame. With all optimizations (downsampled features, fused kernel), this adds 1.8 ms, yielding 127 FPS—still real-time but lower than static mode. For video streams, temporal caching (85% hit rate on sequences with similar consecutive frames) reduces amortized overhead to 0.4 ms per frame.
Table VI provides a detailed runtime breakdown. Table II reports static deployment FPS, which represents the primary use case for industrial applications. Online deployment results are provided in Appendix H.
| Component | Calib. (ms) | Infer. (ms) |
| Backbone + Neck | 4.2 | 4.2 |
| Morphological Analysis | 1.8 | – (precomp.) |
| Bit Allocation (MLP) | 0.4 | – (precomp.) |
| Bit-map Lookup + Quant | – | 0.3 |
| Detection Head | 1.5 | 1.5 |
| Core Pipeline | 7.9 | 6.0 |
| Data I/O + Pre/Post + NMS | – | 0.6 |
| End-to-End Total | – | 6.6 |
| Throughput | – | 151 FPS |
Comparison to Uniform-4bit (156 FPS): The 5 FPS difference (0.3 ms) is due to the per-tile bit-width lookup and non-uniform memory access patterns in the quantization kernel, not the morphological analysis (which runs only at calibration).
V-G Convergence Behavior
Figure 4 compares training dynamics with and without curriculum learning.
@@ V-G Convergence Behavior
Curriculum learning accelerates convergence by gradually exposing the model to more difficult samples while annealing the bit-allocation temperature. With curriculum learning, MCAQ-YOLO reaches 80% mAP at 20k iterations compared to 50k iterations without curriculum ( speedup). The validation mAP variance is also reduced by approximately 60%, indicating more stable optimization dynamics. The curriculum-based strategy reaches 95% of its final performance (81.9% mAP) at 20k iterations, while the non-curriculum variant requires 50k iterations to reach a comparable level (81.1% mAP). The curriculum approach also achieves 3.2 percentage points higher final accuracy (85.6% vs 82.4%), showing improvements in both training efficiency and model performance.
V-H Qualitative Analysis
Figure 5 visualizes the spatial behavior of MCAQ-YOLO on representative examples.
The complexity heatmap highlights boundary- and texture-rich regions (e.g., object contours, fine-grained patterns on clothing), and the resulting bit allocation assigns higher precision to those areas while maintaining lower precision for backgrounds. Notably, the bit allocation varies within objects: more bits are allocated to semantically important details (face, hands, equipment boundaries) than to uniform regions (solid-colored clothing). This adaptive allocation preserves detection accuracy for challenging cases while maximizing compression for simple regions.
@@ VI-A Key Findings and Implications
Curriculum learning is demonstrated to be an effective mechanism for stabilizing quantization-aware training. By gradually increasing the complexity of training samples and adjusting the temperature of bit allocation, faster convergence, lower gradient variance, and improved final accuracy are achieved.